# Understanding Ethical Practices in AI: Insights from a Cross-Role, Cross-Region Survey of AI Development Teams

### v1

@@ 2.2. Ethics in AI Development
Several studies have examined gaps in the practices and organizational structures of AI development teams when implementing ethics principles (Sanderson et al., 2023; Vakkuri et al., 2022, 2019b; Agbese et al., 2023). Hartikainen et al.(Hartikainen et al., 2022) found that developers are often responsible for the early design of AI systems, where they follow flexible practices rather than strict design processes, testing with clients but not end-users. Stahl et al. (Stahl et al., 2022) emphasized the need for organizational awareness and stakeholder communication to implement AI ethics principles. Kawakami et al.’s (Kawakami et al., 2024) study on how power dynamics influence decision-making in AI design highlights the importance of training leadership to effectively communicate and engage diverse stakeholders in ethical decision-making. A study of software developers from Olsen et al. (Olson et al., 2025) found that marginalized demographics report ethical concerns at a higher frequency, and are more aware and sensitive to ethical implementations in software development. However, these groups lack organizational support to address their moral concerns (e.g., 25% of participants from marginalized groups report they did not feel empowered to intervene when ethical issues arise).
@@
These works highlight AI developers’ perceptions and practices in implementing ethics principles into AI systems. This paper builds on previous efforts and is the first to examine AI ethics practices, perceptions, knowledge, and collaboration through the lens of specific roles such as developers, managers, ethicists, and testers within AI development teams. Previous research has typically focused on a small number of developers in a limited set of countries and addressed only a subset of AI ethics principles. In contrast, our study examines various roles involved in AI development, drawing on responses from more than 400 participants in more than 40 countries and covering a broad spectrum of ethics principles. Furthermore, we investigate how factors such as location, gender, company size, and company type influence AI teams’ perceptions, ethical practices, and mitigation strategies, both overall and in relation to individual principles. We also explore participants’ knowledge, perceptions, and practices regarding emerging AI legislation and the potential influence of such regulations on the future development of AI ethics - a topic largely overlooked in prior research.
3. Study Design and Analysis
@@
3.1. Survey Design
3.1.1. Survey Overview
We conducted two large-scale studies, using Qualtrics: one on the Prolific platform (Palan and Schitter, 2018; Tahaei and Vaniea, 2022) and another one on various forums, such as AI-related Sub-Reddits and Quora, X, LinkedIn, HuggingFace, and Kaggle, with AI development team members in 43 countries. Our studies were completed between August and September 2024.
@@ 3.1.1. Survey Overview
We used Prolific for the following reasons: it maintains a pool of active participants who are regularly screened and vetted by the platform, provides a higher pay rate, allows a selection of a more specific pool of participants with basic programming skills and AI familiarity, and is recommended by prior work (Tahaei and Vaniea, 2022; Kaur et al., 2022).
3.1.2. Recruitment Process and Filtering
@@
The sample size was determined in accordance with Prolific’s guidelines and informed by prior similar research (Prybylo et al., 2024), which recommends a minimum of 300 representative participants. We initially pre-screened the Prolific participants based on the following requirements: (a) being at least 18 years old, (b) being fluent in English, (c) using AI tools at least once a week, (d) possessing knowledge of common programming languages, (e) having more than two years of work experience, and (f) working in software-related industries. We used their industry, rather than their professional role, as a filter since Prolific does not allow a role-based selection. We paid an average of \$12.58 per hour to those who completed the survey, which is above the minimum recommended hourly rate by Prolific (\$8 per hour). Prybylo et al. (Prybylo et al., 2024) discuss that although pre-screening via programming questions, such as those recommended by Danilova et al. (Danilova et al., 2021), has its advantages (Tahaei and Vaniea, 2022; Kaur et al., 2022; Danilova et al., 2021; Serafini et al., 2023), it poses several risks, such as overusing the same questions in multiple surveys, their limited scope to only developers and not other roles (e.g., testers or privacy experts), and the accessibility of AI tools like ChatGPT (Brown et al., 2020), which could answer those questions with high accuracy (Danilova et al., 2021). We discuss how we mitigated the pre-screening challenges below and in Section 4. After the initial pre-screening, we recruited 450 participants across the EU, the US, and other countries (the so-called “world pool”) on Prolific. Of those who started the survey, 120 did not finish it and were thus excluded from our analysis.
@@
For the forums, we carefully selected forums that are related to AI development, such as Quora, HuggingFace, and Kaggle, and subreddits, such as (r/LanguageTechnology, r/PromptEngineering, r/MachineLearningJobs, and r/ComputerVision). We also posted the survey on our LinkedIn and academic X profiles, which mainly consisted of computer science researchers and developers. We recruited participants who were (a) at least 18 years old and (b) had at least two years of AI development or research experience. The participants could enter a raffle to obtain one of the four \$25 Amazon gift cards (as per the compensation described in the consent form). We received a total of 356 responses, and excluded those who did not provide consent (as required by our IRB policy) (2) or did not complete the survey (154). Overall, we had 200 complete responses from the forums.
@@
We then conducted another filtering process to remove unreliable responses from both sources. We first removed those participants who completed the survey in under three minutes (twelve from online forums and Prolific), had duplicate answers from other participants (25 from online forums), and those whose written responses were not in English (nine from online forums). We also attempted to remove AI-generated answers from both sources. We identified such responses based on prior research on analyzing differences in linguistic patterns between human and LLM-generated text (Tang et al., 2024). Munoz et al. (Muñoz-Ortiz et al., 2024) found that LLMs tend to have a similar length distribution for similar tasks with a lower lexical diversity than human responses. Their findings are complemented by Guo et al. (Guo et al., 2023), who found that LLMs tend to have longer answers with a less diverse vocabulary in Q&A tasks than humans. We flagged and removed participants with abnormally long answers and those with multiple answers of similar length, vocabulary, grammatical structure, and syntactical patterns that appeared to be AI-generated (i.e., nine participants from Prolific and 15 from online forums). Ultimately, the final participant pool consisted of 276 individuals from Prolific and 138 from online forums.
Table 1. Breakdown of Participants’ Roles by Track
@@
RA: Requirement Analyst, Product Manager 54 (13.0%) 15 39
@@
ISec: Information Security/Privacy Experts 19 (4.6%) 6 13
AE: AI Ethicist, Legal Teams 5 (1.2%) 2 3
@@
3.1.3. Survey Questions
Our survey features a mix of demographic, perception, expertise, and practice questions, all designed in alignment with our RQs (see Section 1). It addresses the challenges identified in prior research, such as the implementation of AI ethics principles (Morley et al., 2023; Rakova et al., 2021; Vakkuri et al., 2022; Griffin et al., 2024), familiarity with AI regulations and their impact (Pant et al., 2024), the organization of AI development teams (Morley et al., 2023; Rakova et al., 2021; Vakkuri et al., 2022; Griffin et al., 2024; Agbese et al., 2023), and team diversity and stakeholder involvement in decision making (Hartikainen et al., 2022; Stahl et al., 2022; Sanderson et al., 2023).
@@ 3.1.3. Survey Questions
All participants were required to provide informed consent in accordance with our university’s IRB policy before proceeding to the main survey. After consenting, we asked all participants the same set of twelve questions, partly related to demographics and their professional experience with AI development. We then divided the remaining questions into two groups: one for those who did not have professional experience developing AI systems (Group A - No AI Development Experience) and another for those with experience (Group B - With AI Development Experience), based on the reponses to questions P11 and P12. We further divided the Group B participants into seven subgroups, based on the roles defined in Table 1, except for the “Others” category, and asked role-specific questions to ensure the survey was not too long. Previous research (Prybylo et al., 2024) shows that participants may not finish a survey if it is longer than 30 minutes. For example, we asked the ISec roles about addressing ethical concerns regarding data privacy and security (B.13.5) and the AI researchers about their challenges for writing ethics statements (B.16.5). Ten participants who chose “Other, please specify” could not be attributed to any roles. They were thus assigned to Group A. Our breakdown loosely follows the software development lifecycle (SDLC) phases. However, we separated the Information Security/Privacy (ISec) and AI Researcher and Ethicist roles to evaluate their perspectives on AI ethics. The complete list of questions (except questions 1-3, which are the required Prolific identification questions and our consent form) can be found here. [^1] Survey questions: https://tinyurl.com/232zaruh
3.2. Study Analysis Process
@@
Our survey results are organized around our RQs, focusing on how various roles perceive, comprehend, and implement AI ethics throughout the SDLC. We also consider the demographic breakdown to provide additional context and allow a more nuanced understanding of the data. Our analysis follows a mixed-method approach, encompassing both quantitative and qualitative methodologies.
3.2.1. Qualitative Analysis
@@
We evaluate the descriptive and open-ended questions through open coding procedures and iterative processes. For questions related to The perceived risks and mitigation strategies associated with each AI ethics principle, the first and second authors conducted a thematic analysis to find recurring patterns and themes. They independently assigned categories for the first 100 responses, then met to discuss their results, resolve the discrepancies, and create a guideline (see Appendix A). For participants’ definitions of AI within their industries, however, the first author used the NIST Taxonomy of Human-AI Use Cases (Theofanos et al., 2023) to categorize their responses, and the last author reviewed the overall results. For the remainder of the open-ended questions, the first author conducted a thematic analysis for individual questions. For all open-ended questions, the last author examined these results to resolve inaccuracies.
3.2.2. Quantitative Analysis
@@
We conducted a Kruskal-Wallis test (Kruskal and Wallis, 1952) to analyze responses to the Likert scale questions and a Chi-Square test (Greenwood and Nikulin, 1996) to assess the association between familiarity with ethical principles and behaviors across various team roles. For perception and experience-based questions, we hypothesize from our RQs that factors such as company size, team roles, geographic location, and gender may influence individuals’ perceptions, practices, and familiarity with AI ethics principles, regulations, and initiatives, ultimately shaping development team practices. To control for Type I errors and reduce the likelihood of false positives, we applied the Bonferroni correction (Shaffer, 1995). Since the Bonferroni correction is very conservative and might increase Type II errors, we discussed the results with respect to $alpha=0.05$ and the adjusted value (i.e., $\frac{0.05}{16}=0.0031$, for our 16 statistical tests).
4. Ethics & Limitations
Ethical Considerations
@@ 4. Ethics & Limitations
This research adheres to our university’s ethical guidelines and was approved by our Institutional Review Board (IRB). All participants agreed to a thorough consent form, including information about the investigators, risks, benefits, compensation, and confidentiality. All participants were informed about their voluntary participation, emphasizing their right to withdraw at any time. No personally identifiable information was collected, and measures were in place to ensure the anonymity, confidentiality, and security of responses. The contact information of all investigators and the IRB team was also included. No participants contacted the investigators or the IRB about the study or the compensation.
Limitations
Similar to any survey study, our analysis relies on participants’ self-reported data. Thus, it may be affected by self-report bias, recall bias, and/or social desirability bias. During the consent process, we informed the participants that the survey pertained to AI practices, but we intentionally did not mention ethics to minimize the impact of self-selection biases. The Prolific user base may not fully represent the diverse population of AI development teams; thus, the survey may face recruitment bias. We diversified our sample by including participants from various AI-related forums alongside those from Prolific. We used multiple screening questions to ensure that recruited participants have experience in AI development activities (Section 3.1.2). We adopted a conservative process to remove participants who appeared to provide AI-generated answers by carefully scrutinizing the write-in responses and investigating semantic qualities of LLM vs. human responses (Tang et al., 2024; Muñoz-Ortiz et al., 2024; Guo et al., 2023). However, AI-generated responses might have affected the study’s outcomes, and we might have removed some professionals. We carefully framed our questions so as not to prompt biased responses and began the survey with questions about AI practices in general to minimize the framing bias. Although we aimed to minimize the risks in answering follow-up write-in questions, there might still be some biases in those questions. We employed statistical analyses to ensure the broad applicability of our findings. To control for Type I errors in the presence of multiple hypothesis tests, we reported our results after employing the Bonferroni correction.
5. Findings
5.1. Demographics

### v2

@@ 2.2 Ethics in AI Development
Other studies have revealed specific challenges associated with implementing AI ethics principles. Khan et al. [khan2023aiethics] conducted a user study with developers and law practitioners, revealing that the biggest challenges in implementing AI ethics principles are a lack of ethical knowledge and unclear regulations. Hartikainen et al.[hartikainen2022human] found that developers are often responsible for the early design of AI systems, where they follow flexible practices rather than strict design processes, testing with clients but not end-users. Stahl et al. [stahl2022organisational] emphasized the need for organizational awareness and stakeholder communication to implement AI ethics principles. Kawakami et al.’s [kawakami2024studyingup] study on how power dynamics influence decision-making in AI design highlights the importance of training leadership to effectively communicate and engage diverse stakeholders in ethical decision-making. A study by Olson et al. [olson2025speaks] shows that marginalized demographics report ethical concerns at a higher frequency, and are more aware and sensitive to ethical implementations in software development. Reuel et al. propose a formal maturity model to measure the organizational and operational maturity of Responsible AI (RAI) [reuel2025responsibleai] and highlight that companies may appear to be more prepared for RAI in terms of organizational structure, internal processes, and frameworks than their practices reveal. For many organizations, RAI was seen as a means of improving revenue, brand performance, or technical performance.
@@
These works highlight AI developers’ perceptions and practices in implementing ethics principles into AI systems. This paper builds on previous efforts and is the first to examine AI ethics practices, perceptions, knowledge, and collaboration through the lens of specific roles such as developers, managers, ethicists, and testers within AI development teams. Previous research has typically focused on a small number of developers in a limited set of countries and addressed only a subset of AI ethics principles. In contrast, our study examines various roles involved in AI development, drawing on responses from over 400 participants in more than 40 countries and covering a broad spectrum of ethics principles. We investigate how factors such as location, gender, company size, and company type influence AI teams’ perceptions, ethics practices, and mitigation strategies, both overall and in relation to individual principles. We also explore participants’ knowledge, perceptions, and practices regarding emerging AI legislation and the potential influence of such regulations on the future development of AI ethics - a topic largely overlooked in prior research.
3 Study Design and Analysis
@@
3.1 Survey Design
3.1.1 Overview
@@
We conducted two large-scale studies, using Qualtrics: one on the Prolific platform [palan2018prolific, Tahaei-CHI2022] and another on various forums, such as AI-related subreddits and Quora, X, LinkedIn, HuggingFace, and Kaggle, with AI development team members in 43 countries. Our studies were completed between August and September 2024. We used Prolific for the following reasons: it maintains a pool of active participants who are regularly screened and vetted by the platform, provides a higher pay rate, allows a selection of a more specific pool of participants with basic programming skills and AI familiarity, and is recommended by prior work [Tahaei-CHI2022, kaur2022recruit].
3.1.2 Recruitment Process and Filtering
@@
The sample size was determined in accordance with Prolific’s guidelines and informed by prior similar research [prybylo2024evaluating], which recommends a minimum of 300 representative participants. We initially pre-screened the Prolific participants based on the following requirements: (a) being at least 18 years old, (b) being fluent in English, (c) using AI tools at least once a week, (d) possessing knowledge of common programming languages, (e) having more than two years of work experience, and (f) working in software-related industries. We used their industry, rather than their professional role, as a filter since Prolific does not allow a role-based selection. We paid an average of \$12.58 per hour to those who completed the survey - above the minimum recommended hourly rate by Prolific (\$8 per hour). Prybylo et al. [prybylo2024evaluating] discuss that although pre-screening via programming questions, such as those recommended by Danilova et al. [danilova2021you], has its advantages [Tahaei-CHI2022, kaur2022recruit, danilova2021you, serafini2023recruitment], it poses several risks, such as overusing the same questions in multiple surveys, their limited scope to only developers and not other roles (e.g., testers or privacy experts), and the accessibility of AI tools like ChatGPT [brown2020language], which could answer those questions with high accuracy [danilova2021you]. We discuss how we mitigated the pre-screening challenges below and in Section 3.2.3. After the initial pre-screening, we recruited 450 participants across the EU, the US, and other countries (i.e., “world pool”) on Prolific. Of those who started the survey, 120 did not finish it and were thus excluded from our analysis.
@@
For the forums, we carefully selected those related to AI development, such as Quora, HuggingFace, and Kaggle, as well as subreddits, like (r/LanguageTechnology, r/PromptEngineering, r/MachineLearningJobs, and r/ComputerVision). We also posted the survey on our LinkedIn and academic X profiles, which mainly consist of computer science researchers and developers. We recruited participants who were (a) at least 18 years old and (b) had at least two years of AI development or research experience. The participants could enter a raffle to obtain one of the four \$25 Amazon gift cards (as per the compensation described in the consent form). We received a total of 356 responses, and excluded those who did not provide consent (as required by our IRB policy) (2) or did not complete the survey (154). Overall, we received 200 complete responses from the forums.
@@
We then conducted another filtering process to remove unreliable responses from both sources. We first removed those who completed the survey in under three minutes (twelve from online forums and Prolific), had duplicate answers from other participants (25 from online forums), and those whose written responses were not in English (nine from online forums). We also attempted to remove AI-generated answers from both sources. We identified such responses based on prior research on analyzing differences in linguistic patterns between human and LLM-generated text [tang2024science]. Munoz et al. [munoz2024contrasting] found that LLMs tend to have a similar length distribution for similar tasks with a lower lexical diversity than human responses. Their findings are complemented by Guo et al. [guo2023how], who found that LLMs tend to have longer answers with a less diverse vocabulary in Q&A tasks than humans. We flagged and removed participants with abnormally long answers and those with multiple answers of similar length, vocabulary, grammatical structure, and syntactical patterns that appeared to be AI-generated (i.e., nine participants from Prolific and 15 from online forums). Ultimately, the final participant pool consisted of 276 individuals from Prolific and 138 from online forums.
Table 1: Breakdown of Participants’ Roles
@@
RA: Requirements Analyst 54 (13.0%) 15 39
@@
ISec: Information Security/Privacy Expert 19 (4.6%) 6 13
AE: AI Ethicist, Legal Team 5 (1.2%) 2 3
@@
3.1.3 Survey Questions
Our survey features a mix of demographic, perception, expertise, and practice questions, all designed in alignment with our RQs (see Section 1). The survey design builds upon the methodological framework established by Prybylo et al., [prybylo2024evaluating], extending their role and demographic-based analysis of privacy perceptions and practices to the broader domain of AI ethics. Following this precedent, we utilized Likert-scale items to quantify perceived impact, familiarity, and frequency of ethical practices, allowing for robust statistical comparisons across various demographics. We also leveraged the findings, challenges and the gaps identified in prior research, such as the implementation of AI ethics principles [morley2023operationalising, rakova2021where, vakkuri2022how, griffin2024ethical], familiarity with AI regulations and their impact [pant2024ethics], the organization of AI development teams and culture [morley2023operationalising, rakova2021where, vakkuri2022how, griffin2024ethical, agbese2023implementing], and team diversity and stakeholder involvement in decision making [hartikainen2022human, stahl2022organisational, sanderson2023ai]to formulate the questions. Although we did not conduct a pilot study, we asked members of our research lab to complete the survey, and iteratively refined the questions based on their feedback and responses. Lastly, as one of the authors teaches computer and business ethics courses, the survey questions were informed based on his experience and interaction with both computer science and business students in his classes.
@@ 3.1.3 Survey Questions
All participants were required to provide informed consent in accordance with our university’s IRB policy before proceeding to the main survey. After obtaining consent, we asked all participants the same set of twelve questions, partly related to demographics and their professional experience with AI development. We then divided the remaining questions into two groups: one for those who did not have professional experience developing AI systems (Group A - No AI Development Experience) and another for those with experience (Group B - With AI Development Experience), based on the responses to questions P11 and P12. We further divided the Group B participants into seven subgroups, based on the roles defined in Table 1, except for the “Others” category, and asked role-specific questions to ensure the survey was not too long. Previous research [prybylo2024evaluating] shows that participants may not finish a survey if it is longer than 30 minutes. For example, we asked the ISec roles about addressing ethical concerns regarding data privacy and security (B.13.5) and the AI researchers about their challenges in writing ethics statements (B.16.5). Ten participants who chose “Other, please specify” could not be attributed to any specific role and were thus assigned to Group A. Our breakdown loosely follows the software development lifecycle (SDLC) phases. However, we separated the Information Security/Privacy (ISec) and AI Researcher and Ethicist roles to evaluate their perspectives on AI ethics. The complete list of questions (except questions 1-3, which are the required Prolific identification questions and our consent form) can be found here. [^1] Survey questions: https://tinyurl.com/232zaruh
3.2 Study Analysis Process
@@
Our survey results are organized around our RQs, focusing on how various roles perceive, comprehend, and implement AI ethics throughout the SDLC. We also consider the demographic breakdown to provide additional context and allow a more nuanced understanding of the data. Our analysis employs a mixed-method approach, combining both quantitative and qualitative methodologies.
3.2.1 Qualitative Analysis
@@
We evaluated the descriptive and open-ended questions through open coding procedures and iterative processes. For questions related to The perceived risks and mitigation strategies associated with each AI ethics principle, the first and second authors conducted a thematic analysis to find recurring patterns and themes. They independently assigned categories for the first 100 responses, then met to discuss their results, resolve the discrepancies, and create a guideline (see Appendix C). For participants’ definitions of AI within their industries, however, the first author used the NIST Taxonomy of Human-AI Use Cases [nist2024taxonomy] to categorize their responses, and the last author reviewed the overall results. For the remainder of the open-ended questions, the first author conducted a thematic analysis for individual questions. For all open-ended questions, the last author (who is an ethics researcher) examined these results to resolve inaccuracies.
3.2.2 Quantitative Analysis
@@
We conducted a Kruskal-Wallis test [kruskal1952use] to analyze responses to the Likert scale questions and a Chi-Square test [greenwood1996guide] for the association between AI ethics principles or AI governance initiatives familiarity and various demographic factors. For perception and experience-based questions, we hypothesize from our RQs that factors such as company size, team roles, location, and gender may influence individuals’ perceptions, practices, and familiarity with AI ethics principles and governance initiatives, ultimately shaping development team practices. To control for Type I errors and reduce the likelihood of false positives, we applied the Bonferroni correction [Bonferroni]. Since the Bonferroni correction is very conservative and may increase Type II errors, we discussed the results with respect to $alpha=0.05$ and the adjusted value (i.e., $\frac{0.05}{16}=0.0031$).
We conducted least absolute shrinkage and selection operator (LASSO) [tibshirani1996lasso] regression to identify which demographics, practices, and perceptions are most associated with: (1) familiarity with AI ethics principles and governance initiatives, (2) frequency of considering AI ethics principles, and (3) perceived impact of AI governance initiatives. We used the Likert-scale responses as continuous dependent variables in all models. For potential explanatory variables, we used the questions listed in Table 31.
We preprocessed features according to their measurement type: ordinal variables (age groups, company size, experience levels, and frequency scales) were coded numerically to preserve rank order; categorical variables (education level, company type, role) were one-hot encoded; multi-select questions (AI applications, tools used, mitigation strategies) were converted to binary indicator variables; and binary variables (e.g., company builds AI) were coded as 0/1. This step yielded 118 and 179 features for Groups A and B, respectively. Following best practices [hastie2009elements], we employed a two-stage procedure: first, we conducted a LASSO analysis with 5-fold cross-validation using the one-standard-error rule to select important predictors; second, we re-estimated coefficients using standard linear regression on only the selected features to obtain unbiased estimates. We reported post-LASSO standardized coefficients, which represent the change in the outcome variable (on the original Likert scale) associated with a one-standard-deviation change in each predictor. We evaluated the model performance using R² (proportion of variance explained) on a held-out test set (20% of data).
Finally, we employed linear mixed-effects models to analyze our survey data, taking into account the hierarchical structure in which practitioners are nested within professional roles. This approach is appropriate because individuals within the same role may share similar organizational contexts and professional norms [stroup2012glmm].
We specify two models to test our hypotheses about factors influencing awareness of AI ethics principles (RQ2) and how they are considered in practice (RQ3).
Model 1 (Principle Familiarity) predicts practitioners’ mean familiarity with nine AI ethics principles (scale 1–5) based on years of AI development experience (ordinal 1–4 representing 1–2 years through 10+ years), active AI development involvement (binary), and mean familiarity with eight AI governance initiatives (scale 1–5), with random intercepts for professional roles. Model 2 (Ethics Consideration Frequency) predicts how often practitioners consider ethics in their work (scale 1–5) based on mean familiarity with AI ethics principles, mean familiarity with AI governance initiatives, and years of experience, with random intercepts for professional role. Professional roles included all roles as shown in Table 1.
Prior to interpretation, we performed diagnostics to assess normality assumptions. Outcome Variable Distributions: We examined distributions of all outcome variables prior to modeling. The familiarity principle shows left skewness (skewness = 0.790-0.790), as does the ethics consideration frequency (skewness = 0.731-0.731). Familiarity with AI governance initiatives is approximately symmetric (skewness = 0.152-0.152). Shapiro-Wilk tests rejected the normality assumption for all three outcomes ($p<0.001$). Model residuals show some deviation from perfect normality: Shapiro-Wilk tests yielded $W=0.974$ ($p=0.0001$) for Model 1 and $W=0.983$ ($p=0.003$) for Model 2. However, we consider these models acceptable given the mild nature of departures (W-statistics near 0.98), large sample sizes, and significant effects ($p<0.01$). Box-Cox Transformation Exploration: To address non-normality, we used Box-Cox power transformations for each outcome variable. For principle familiarity, the optimal $\lambda=1.99$ reduced skewness from 0.790 to 0.167, though normality tests remained significant ($W=0.972$, $p<0.001$). For frequency of considering AI ethics principles, $\lambda=1.86$ reduced skewness from 0.731 to 0.195 ($W=0.873$, $p<0.001$). For AI governance initiative familiarity, $\lambda=0.90$ provided minimal improvement, with skewness increasing slightly from 0.152 to 0.202 ($W=0.937$, $p<0.001$). While Box-Cox transformations reduced skewness for two outcomes, they did not achieve normality; therefore, we did not use these transformed data in our analysis.
3.2.3 Ethical Considerations
@@
This research adheres to our university’s ethical guidelines and was approved by our Institutional Review Board (IRB). All participants agreed to a thorough consent form that included information about the investigators, risks, benefits, compensation, and confidentiality. All participants were informed about their voluntary participation, with the emphasis on their right to withdraw at any time. No personally identifiable information was collected, and measures were in place to ensure the anonymity, confidentiality, and security of responses. The contact information of all investigators and the IRB team was also included. No participants contacted the investigators or the IRB about the study or the compensation.
4 Findings
4.1 Demographics